# It’s TIME: Towards the Next Generation of Time Series Forecasting Benchmarks

### v1

@@ 1 Introduction
  • We evaluate 12 representative TSFMs on TIME and develop an interactive leaderboard supporting multi-granular analysis, combining quantitative results with qualitative visualization to improve the interpretability and actionability of benchmark outcomes.
@@ 5.1 Structural tsfeatures for Pattern Definition
Figure 4: Empirical density distributions of our structural tsfeatures ($F_{1}$$F_{6}$) across the benchmark variates. Stationarity ($F_{7}$) is omitted here as it is a binary indicator, with non-stationary variates accounting for 7.6% of the total.
@@ Results.
Figure 5 illustrates the overall performance on point and probabilistic forecasting across all tasks. TimesFM-2.5 achieves the lowest MASE, while Chronos2 leads in probabilistic forecasting with the best CRPS. A distinct performance progression is observed where newer iterations (e.g., TimesFM-2.5 vs. 2.0/1.0, and Moirai2 vs. Moirai, Chronos2 vs. Chronos-bolt) consistently outperform their predecessors. This confirms that recent advancements in TSFMs represent genuine capability improvements rather than overfitting to the data bias in established benchmarks. Analyzing the profile of the top-5 performing models reveals distinct design trends: decoder-only frameworks (TimesFM-2.5, ToTo, Moirai2) and quantile-based mechanisms are widely adopted.Notably, the leading positions are generally occupied by the most recently released models. This trend not only highlights the development of the TSFM field but also validates the rationality of our benchmark in accurately capturing these genuine advancements.
@@ 6.3 Pattern-specific Performance
In terms of Seasonality Correlation ($F_{4}$), the evolution of TSFMs brings performance gains across both stable and unstable seasonal patterns. For early/weaker models (Moirai, VisionTS++, and TimesFM-1.0), the modeling effectiveness on stable periods is relatively close to that of S-Naive, which results in a significant gap compared to the much larger gains seen on unstable sequences. With further development, this gap has narrowed significantly. Most recent TSFMs show a relatively small performance difference between stable and unstable sequences. For leading models like TimesFM-2.5 and Chronos2, the performance gains on stable sequences actually surpass those on unstable ones. This suggests that recent TSFMs have become more robust to seasonality stability and provide consistent improvements.
@@
For the majority of TSFMs, the relative gains are more pronounced on non-stationary sequences. This trend highlights the limitations of S-Naive, which struggle to adapt to the shifting statistical properties of non-stationary data. TSFMs, however, demonstrate a robust capacity to model these dynamic changes, resulting in a larger performance gap over the baseline in non-stationary scenarios. Notably, the model rankings exhibit high sensitivity to stationarity. For instance, ToTo ranks second on non-stationary data but drops to fifth on stationary sequences. Conversely, Chronos2 dominates stationary sequences as the top performer but falls to sixth place when dealing with non-stationary data. This shift suggests that different TSFMs possess unique inductive biases, with some optimized for handling non-stationary shifts and others prioritizing precision in stationary dynamics.
@@
Although model rankings exhibit specific variations across different pattern variates, the overall profile remains largely consistent. In most cases, TimesFM-2.5 and Chronos-2 maintain their status as the top two performers, while Moirai2, TiRex and ToTo typically fluctuate within the third to fifth positions. Notably, we find that these feature-specific rankings do not align with the overall rankings presented in Section 6.2. While task-level aggregation is the common practice in existing benchmarks (Aksu et al., 2024; Shchur et al., 2025), our first attempt at variate-level aggregation reveals that rankings are significantly influenced by the level of aggregation. Determining the most appropriate aggregation level remains an open question, highlighting that a single leaderboard ranking is not the sole metric for success and requires more granular quantitative analysis.
@@ C.2 Models
Table 3: Model configurations used in evaluation. Context length refers to the maximum number of historical time steps used as input. Input mode indicates whether the model natively supports multivariate time series or processes each variate independently (univariate). Output type specifies whether models produce quantile forecasts or sample-based probabilistic forecasts.
@@
Chronos2 amazon/chronos-2 4000 Multivariate Quantile
@@
Toto Datadog/Toto-Open-Base-1.0 4096 Multivariate Sample (100)
Sundial thuml/sundial-base-128m 2880 Univariate Sample (100)
@@
Chronos-bolt amazon/chronos-bolt-base 4000 Univariate Quantile
Moirai Salesforce/moirai-1.1-R-base 4000 Multivariate Sample (100)
@@
Models with quantile output produce 9 quantile forecasts at levels $\{0.1,0.2,\ldots,0.9\}$. Models with sample-based output generate 100 Monte Carlo samples for probabilistic forecasting.
@@
Chronos2, VisionTS++, Toto, and Moirai (1.1) natively support multivariate time series input. For univariate models, we flatten multivariate series into independent univariate sequences and forecast each variate separately. For Moirai (1.1), when GPU memory is insufficient for multivariate inference, the model automatically falls back to univariate mode to complete the evaluation.
@@ Appendix E Full Results
Metric TimesFM 2.5 Chronos2 Kairos Moirai 2.0 VisionTS++ TiRex Toto Sundial TimesFM 2.0 Chronos-bolt Moirai TimesFM 1.0
MASE 0.667 0.712 0.738 0.701 0.803 0.697 0.696 0.755 0.717 0.784 0.823 0.787
CRPS 0.559 0.551 0.635 0.580 0.664 0.581 0.576 0.654 0.613 0.621 0.657 0.679
MASE Rank 3.10 4.58 7.86 5.86 9.57 4.72 5.76 8.32 6.88 9.18 10.85 10.45
CRPS Rank 3.43 2.67 9.37 5.71 9.30 4.65 5.43 10.74 7.55 8.27 9.87 11.34
@@
Trend $F_{k}=1$ MASE 0.601 0.618 0.678 0.633 0.725 0.642 0.651 0.66 0.653 0.707 0.788 0.759
CRPS 0.526 0.543 0.614 0.558 0.642 0.568 0.57 0.606 0.591 0.632 0.691 0.698
$F_{k}=0$ MASE 0.651 0.639 0.724 0.665 0.767 0.67 0.669 0.697 0.677 0.723 0.769 0.734
CRPS 0.579 0.563 0.68 0.594 0.683 0.587 0.589 0.652 0.625 0.652 0.685 0.699
T. Linearity $F_{k}=1$ MASE 0.613 0.613 0.677 0.633 0.726 0.645 0.666 0.664 0.649 0.696 0.783 0.741
CRPS 0.542 0.542 0.622 0.563 0.646 0.568 0.586 0.616 0.593 0.626 0.694 0.692
$F_{k}=0$ MASE 0.632 0.642 0.721 0.661 0.761 0.664 0.651 0.687 0.679 0.733 0.776 0.756
CRPS 0.555 0.563 0.663 0.584 0.673 0.584 0.569 0.636 0.618 0.656 0.682 0.704
Seasonality $F_{k}=1$ MASE 0.586 0.593 0.694 0.619 0.772 0.632 0.648 0.653 0.64 0.715 0.843 0.766
CRPS 0.529 0.537 0.654 0.563 0.701 0.568 0.581 0.614 0.591 0.655 0.758 0.731
$F_{k}=0$ MASE 0.662 0.664 0.702 0.676 0.713 0.679 0.67 0.7 0.689 0.713 0.717 0.73
CRPS 0.569 0.568 0.628 0.584 0.618 0.584 0.575 0.638 0.62 0.625 0.621 0.665
S. Correlation $F_{k}=1$ MASE 0.617 0.619 0.701 0.645 0.77 0.662 0.669 0.676 0.665 0.718 0.853 0.783
CRPS 0.545 0.551 0.646 0.577 0.684 0.587 0.59 0.625 0.608 0.644 0.751 0.734
$F_{k}=0$ MASE 0.627 0.634 0.695 0.647 0.718 0.646 0.649 0.674 0.661 0.709 0.714 0.716
CRPS 0.552 0.553 0.637 0.57 0.636 0.565 0.567 0.626 0.603 0.637 0.633 0.665
R. ACF1 $F_{k}=1$ MASE 0.603 0.619 0.681 0.633 0.73 0.642 0.656 0.666 0.656 0.708 0.799 0.767
CRPS 0.533 0.549 0.625 0.563 0.651 0.569 0.58 0.619 0.601 0.636 0.709 0.716
$F_{k}=0$ MASE 0.649 0.638 0.72 0.664 0.76 0.671 0.662 0.687 0.673 0.721 0.754 0.724
CRPS 0.57 0.556 0.663 0.586 0.67 0.584 0.576 0.634 0.611 0.646 0.662 0.675
Stationarity $F_{k}=1$ MASE 0.626 0.624 0.699 0.65 0.74 0.658 0.665 0.672 0.665 0.713 0.788 0.755
CRPS 0.554 0.551 0.645 0.578 0.658 0.58 0.585 0.625 0.61 0.642 0.698 0.708
$F_{k}=0$ MASE 0.583 0.652 0.685 0.613 0.777 0.62 0.6 0.714 0.639 0.721 0.697 0.678
CRPS 0.495 0.558 0.601 0.52 0.676 0.534 0.508 0.636 0.558 0.624 0.596 0.598
Complexity $F_{k}=1$ MASE 0.657 0.651 0.692 0.662 0.692 0.663 0.668 0.687 0.674 0.696 0.695 0.708
CRPS 0.569 0.559 0.626 0.574 0.602 0.569 0.575 0.629 0.61 0.613 0.609 0.653
$F_{k}=0$ MASE 0.596 0.609 0.702 0.634 0.784 0.648 0.652 0.666 0.655 0.728 0.851 0.781
CRPS 0.533 0.547 0.653 0.572 0.707 0.581 0.58 0.623 0.601 0.663 0.756 0.735
@@
Dataset Metric TimesFM 2.5 Chronos2 Kairos Moirai 2.0 VisionTS++ TiRex Toto Sundial TimesFM 2.0 Chronos-bolt Moirai TimesFM 1.0
Water Quality-Darwin MASE (Raw) 0.993 1.089 1.220 1.092 1.270 1.024 1.007 1.292 1.182 1.214 1.428 1.098
@@
MASE (Norm.) 0.739 0.839 0.928 0.816 0.989 0.766 0.754 0.953 0.888 0.932 1.072 0.827
CRPS (Norm.) 0.504 0.466 0.694 0.584 0.795 0.553 0.500 0.745 0.645 0.565 0.675 0.630
MASE (Rank) 1.667 4.667 10.333 4.667 9.667 3 2.333 10.333 8.667 9.333 12 5.333
@@
Current Velocity/5T MASE (Raw) 0.567 0.699 0.790 0.787 0.876 0.639 0.665 0.621 0.608 0.796 0.848 0.719
CRPS (Raw) 0.276 0.283 0.343 0.315 0.293 0.278 0.289 0.291 0.314 0.319 0.289 0.361
MASE (Norm.) 0.545 0.684 0.733 0.715 0.875 0.609 0.624 0.596 0.583 0.767 0.811 0.675
CRPS (Norm.) 0.641 0.657 0.769 0.720 0.683 0.645 0.670 0.678 0.720 0.728 0.672 0.815
MASE (Rank) 1 6.667 10 9 10 4 4 3.667 2.667 9.667 10.667 7.333
CRPS (Rank) 2 3.667 10 8 6.333 2 5.667 8 7 8.333 6.333 11.333
Current Velocity/10T MASE (Raw) 0.729 0.799 0.956 0.929 1.183 0.788 0.833 0.907 0.804 1.065 1.281 0.897
@@
MASE (Norm.) 0.601 0.666 0.787 0.752 1.005 0.647 0.678 0.738 0.662 0.876 1.046 0.734
CRPS (Norm.) 0.649 0.633 0.786 0.757 0.862 0.666 0.691 0.749 0.712 0.779 0.903 0.766
MASE (Rank) 1 4.333 8.667 7.667 11.333 2.333 4.333 6.667 4 10.667 12.333 6.333
@@
Current Velocity/15T MASE (Raw) 0.695 1.330 0.815 0.734 0.884 0.703 0.738 0.768 0.738 1.391 0.904 0.814
@@
MASE (Norm.) 0.648 1.264 0.754 0.682 0.840 0.654 0.685 0.714 0.685 1.319 0.836 0.752
CRPS (Norm.) 0.659 0.655 0.697 0.675 0.708 0.658 0.677 0.710 0.689 0.678 0.716 0.717
MASE (Rank) 1.333 12 8 4 8.667 1.667 4 6 4 13 9.667 7.667
CRPS (Rank) 2.333 2 9 5.667 7.667 2 6 10 6.333 6.333 10.333 10.333
Current Velocity/20T MASE (Raw) 0.762 0.881 0.866 0.860 1.017 0.855 0.814 0.897 0.796 1.168 1.094 0.864
CRPS (Raw) 0.316 0.305 0.350 0.339 0.368 0.391 0.324 0.338 0.329 0.430 0.367 0.355
MASE (Norm.) 0.644 0.761 0.736 0.723 0.898 0.746 0.687 0.758 0.673 1.021 0.927 0.726
CRPS (Norm.) 0.648 0.626 0.716 0.693 0.760 0.796 0.665 0.695 0.673 0.881 0.754 0.722
MASE (Rank) 1 6.667 6.667 5.667 9.333 5 3.667 8.667 2.333 12.333 11.333 6.667
CRPS (Rank) 2 1 8.667 6.333 7.667 6 4.667 7 4.667 10.667 10.333 9
Current Velocity/H MASE (Raw) 1.062 1.303 1.373 1.121 1.482 1.123 1.175 1.775 1.258 1.531 1.405 1.457
@@
MASE (Norm.) 0.650 0.808 0.844 0.684 0.930 0.681 0.715 1.082 0.760 0.945 0.869 0.887
CRPS (Norm.) 0.633 0.623 0.765 0.661 0.748 0.625 0.666 0.677 0.729 0.725 0.729 0.832
MASE (Rank) 1.333 6.333 7.333 2.667 10 2 4 12.333 5.667 11 8 8.667
@@
CPHL/15T MASE (Raw) 0.977 0.966 0.992 0.994 0.973 0.957 0.938 0.922 0.944 0.975 0.998 1.100
CRPS (Raw) 0.213 0.215 0.225 0.216 0.213 0.213 0.208 0.216 0.212 0.217 0.222 0.252
MASE (Norm.) 0.657 0.650 0.666 0.665 0.661 0.645 0.634 0.623 0.637 0.655 0.675 0.735
CRPS (Norm.) 0.615 0.618 0.647 0.620 0.622 0.614 0.601 0.626 0.611 0.624 0.643 0.724
MASE (Rank) 7.667 4.667 10 7.333 7.333 4.333 3.667 1.333 4 6 10 11.667
CRPS (Rank) 5 5 10 5.333 5.333 4.667 3.333 6 4.667 7.333 9.333 12
CPHL/30T MASE (Raw) 1.406 1.383 1.482 1.499 1.490 1.437 1.432 1.451 1.451 1.452 1.583 1.515
CRPS (Raw) 0.248 0.249 0.270 0.267 0.266 0.262 0.255 0.287 0.262 0.265 0.285 0.286
MASE (Norm.) 0.765 0.752 0.805 0.816 0.808 0.781 0.779 0.789 0.790 0.789 0.862 0.822
CRPS (Norm.) 0.689 0.691 0.748 0.740 0.737 0.727 0.709 0.796 0.727 0.733 0.789 0.792
MASE (Rank) 2 1 9 9.667 8 4 3.667 6.333 6.333 6 12 10
CRPS (Rank) 1.333 1.667 8.667 7.667 6.333 4.667 3.333 11.333 5.667 5.667 10.667 11
CPHL/H MASE (Raw) 1.645 1.666 1.699 1.699 1.828 1.718 1.748 1.841 1.705 1.738 1.801 1.869
CRPS (Raw) 0.282 0.292 0.312 0.303 0.315 0.300 0.302 0.351 0.308 0.315 0.312 0.353
MASE (Norm.) 0.800 0.807 0.825 0.824 0.890 0.831 0.841 0.888 0.827 0.843 0.875 0.900
CRPS (Norm.) 0.743 0.766 0.815 0.795 0.829 0.785 0.791 0.916 0.805 0.826 0.820 0.918
MASE (Rank) 1.667 1.667 4.333 5 10 5.333 6 10.333 5.333 7.333 9.667 11.333
CRPS (Rank) 1 2 8 5.333 8.333 3.333 5 11.333 7 8 7.333 11.333
Coastal T-S/5T MASE (Raw) 1.019 1.391 1.135 1.078 1.127 1.051 1.033 1.171 1.071 1.440 1.199 1.139
CRPS (Raw) 0.043 0.042 0.055 0.043 0.046 0.042 0.044 0.057 0.054 0.049 0.045 0.056
MASE (Norm.) 0.733 1.014 0.810 0.774 0.816 0.755 0.743 0.830 0.768 1.048 0.856 0.818
CRPS (Norm.) 0.504 0.489 0.635 0.507 0.534 0.489 0.507 0.669 0.616 0.566 0.527 0.645
MASE (Rank) 1 11.667 7.333 4.667 7.333 3 2 8.333 4.333 12.667 9.333 8

### v2

# arXiv:2602.12147v2[cs.LG] 27 Feb 2026

@@ 1 Introduction
  • We evaluate 12 representative TSFMs on TIME and develop an interactive leaderboard supporting multi-granular analysis, combining quantitative results with qualitative visualization to improve the interpretability and actionability of benchmark outcomes. Codes are available at https://github.com/zqiao11/TIME.
@@ 5.1 Structural tsfeatures for Pattern Definition
Figure 4: Empirical density distributions of our structural tsfeatures ($F_{1}$$F_{6}$) across the benchmark variates. Stationarity ($F_{7}$) is omitted here as it is a binary indicator, with non-stationary variates accounting for 16.6% of the total.
@@ Results.
Figure 5 illustrates the overall performance on point and probabilistic forecasting across all tasks. While using distinct architectures, Chronos-2, TimesFM-2.5, and TiRex emerge as the top-3 performers, consistently achieving the lowest MASE and CRPS scores across the benchmark. A distinct performance progression is observed where newer iterations (e.g., TimesFM-2.5 vs. 2.0/1.0, and Moirai2 vs. Moirai, Chronos-2 vs. Chronos-bolt) consistently outperform their predecessors. This confirms that recent advancements in TSFMs represent genuine capability improvements rather than overfitting to the data bias in established benchmarks. Analyzing the profile of the top-5 performing models reveals distinct design trends: decoder-only frameworks (TimesFM-2.5, Toto, Moirai2) and quantile-based output mechanisms are widely adopted. Notably, the leading positions are generally occupied by the most recently released models. This trend not only highlights the development of the TSFM field but also validates the rationality of our benchmark in accurately capturing these genuine advancements.
@@ 6.3 Pattern-specific Performance
In terms of Seasonality Correlation ($F_{4}$), the evolution of TSFMs brings performance gains across both stable and unstable seasonal patterns. For early/weaker models (Moirai, VisionTS++, and TimesFM-1.0), the modeling effectiveness on stable periods is relatively close to that of S-Naive, which results in a significant gap compared to the much larger gains seen on unstable sequences. With further development, this gap has narrowed significantly. Most recent TSFMs show a relatively small performance difference between stable and unstable sequences. For leading models like Chronos-2 and TimesFM-2.5, the performance gains on stable sequences actually surpass those on unstable ones. This suggests that recent TSFMs have become more robust to seasonality stability and provide consistent improvements.
@@
For the majority of TSFMs, the relative gains are more pronounced on non-stationary sequences. This trend highlights the limitations of S-Naive, which struggle to adapt to the shifting statistical properties of non-stationary data. TSFMs, however, demonstrate a robust capacity to model these dynamic changes, resulting in a larger performance gap over the baseline in non-stationary scenarios. Notably, the model rankings exhibit high sensitivity to stationarity. For instance, Toto ranks third on non-stationary data but drops to sixth on stationary sequences. Conversely, Chronos-2 dominates stationary sequences as the top performer, but is surpassed by TimesFM-2.5 when dealing with non-stationary data. This shift suggests that different TSFMs possess unique inductive biases, with some optimized for handling non-stationary shifts and others prioritizing precision in stationary dynamics.
@@
Although model rankings exhibit specific variations across different pattern variates, the overall profile remains largely consistent. In most cases, Chronos-2 and TimesFM-2.5 maintain their status as the top two performers, while Moirai2, TiRex and Toto typically fluctuate within the third to fifth positions. Notably, we find that these feature-specific rankings do not align with the overall rankings presented in Section 6.2. While task-level aggregation is the common practice in existing benchmarks (Aksu et al., 2024; Shchur et al., 2025), our first attempt at variate-level aggregation reveals that rankings are significantly influenced by the level of aggregation. Determining the most appropriate aggregation level remains an open question, highlighting that a single leaderboard ranking is not the sole metric for success and requires more granular quantitative analysis.
@@ C.2 Models
Table 3: Model configurations used in evaluation. Context length refers to the maximum number of historical time steps used as input. Input mode indicates whether the model natively supports multivariate time series or processes each variate independently (univariate). Output type specifies whether models produce quantile forecasts or distribution-based probabilistic forecasts.
@@
Chronos-2 amazon/chronos-2 8192 Multivariate Quantile
@@
Toto Datadog/Toto-Open-Base-1.0 4096 Multivariate Distribution
Sundial thuml/sundial-base-128m 2880 Univariate Distribution
@@
Chronos-bolt amazon/chronos-bolt-base 2048 Univariate Quantile
Moirai Salesforce/moirai-1.1-R-base 4000 Multivariate Distribution
@@
Models with quantile output produce 9 quantile forecasts at levels $\{0.1,0.2,\ldots,0.9\}$. Models with dstribution-based output generate 100 Monte Carlo samples for probabilistic forecasting.
@@
Chronos-2, VisionTS++, Toto, and Moirai (1.1) natively support multivariate time series input. For univariate models, we flatten multivariate series into independent univariate sequences and forecast each variate separately. For Moirai, when GPU memory is insufficient for multivariate inference, the model automatically falls back to univariate mode to complete the evaluation.
@@ Appendix E Full Results
Metric TimesFM 2.5 Chronos-2 Kairos Moirai 2.0 VisionTS++ TiRex Toto Sundial TimesFM 2.0 Chronos-bolt Moirai TimesFM 1.0
MASE 0.667 0.660 0.738 0.701 0.803 0.681 0.696 0.755 0.717 0.731 0.823 0.787
CRPS 0.559 0.549 0.635 0.580 0.664 0.565 0.576 0.654 0.613 0.612 0.657 0.679
MASE Rank 3.12 2.42 7.58 5.76 8.87 3.99 5.53 7.88 6.66 7.15 10.09 9.79
CRPS Rank 3.22 2.32 8.29 5.31 8.22 3.64 4.98 9.41 6.88 7.35 8.74 9.98
@@
Trend $F_{k}=1$ MASE 0.601 0.588 0.678 0.633 0.725 0.633 0.651 0.66 0.653 0.679 0.788 0.759
CRPS 0.526 0.514 0.614 0.558 0.642 0.554 0.57 0.606 0.591 0.6 0.691 0.698
$F_{k}=0$ MASE 0.651 0.63 0.724 0.665 0.767 0.668 0.669 0.697 0.677 0.714 0.769 0.734
CRPS 0.579 0.554 0.68 0.594 0.683 0.584 0.589 0.652 0.625 0.642 0.685 0.699
T. Linearity $F_{k}=1$ MASE 0.613 0.593 0.677 0.633 0.726 0.64 0.666 0.664 0.649 0.677 0.783 0.741
CRPS 0.542 0.521 0.622 0.563 0.646 0.56 0.586 0.616 0.593 0.605 0.694 0.692
$F_{k}=0$ MASE 0.632 0.619 0.721 0.661 0.761 0.657 0.651 0.687 0.679 0.711 0.776 0.756
CRPS 0.555 0.541 0.663 0.584 0.673 0.574 0.569 0.636 0.618 0.632 0.682 0.704
Seasonality $F_{k}=1$ MASE 0.586 0.564 0.694 0.619 0.772 0.626 0.648 0.653 0.64 0.691 0.843 0.766
CRPS 0.529 0.507 0.654 0.563 0.701 0.559 0.581 0.614 0.591 0.628 0.758 0.731
$F_{k}=0$ MASE 0.662 0.653 0.702 0.676 0.713 0.672 0.67 0.7 0.689 0.697 0.717 0.73
CRPS 0.569 0.558 0.628 0.584 0.618 0.576 0.575 0.638 0.62 0.607 0.621 0.665
S. Correlation $F_{k}=1$ MASE 0.617 0.593 0.701 0.645 0.77 0.657 0.669 0.676 0.665 0.698 0.853 0.783
CRPS 0.545 0.524 0.646 0.577 0.684 0.579 0.59 0.625 0.608 0.623 0.751 0.734
$F_{k}=0$ MASE 0.627 0.618 0.695 0.647 0.718 0.64 0.649 0.674 0.661 0.689 0.714 0.716
CRPS 0.552 0.537 0.637 0.57 0.636 0.556 0.567 0.626 0.603 0.613 0.633 0.665
R. ACF1 $F_{k}=1$ MASE 0.603 0.585 0.681 0.633 0.73 0.635 0.656 0.666 0.656 0.681 0.799 0.767
CRPS 0.533 0.516 0.625 0.563 0.651 0.561 0.58 0.619 0.601 0.607 0.709 0.716
$F_{k}=0$ MASE 0.649 0.633 0.72 0.664 0.76 0.665 0.662 0.687 0.673 0.71 0.754 0.724
CRPS 0.57 0.551 0.663 0.586 0.67 0.575 0.576 0.634 0.611 0.632 0.662 0.675
Stationarity $F_{k}=1$ MASE 0.625 0.604 0.694 0.647 0.738 0.649 0.667 0.671 0.661 0.694 0.789 0.743
CRPS 0.548 0.527 0.635 0.571 0.652 0.564 0.583 0.619 0.6 0.615 0.694 0.691
$F_{k}=0$ MASE 0.609 0.613 0.714 0.643 0.766 0.644 0.622 0.696 0.674 0.692 0.736 0.773
CRPS 0.548 0.548 0.671 0.583 0.696 0.582 0.558 0.658 0.63 0.631 0.66 0.732
Complexity $F_{k}=1$ MASE 0.657 0.647 0.692 0.662 0.692 0.661 0.668 0.687 0.674 0.69 0.695 0.708
CRPS 0.569 0.555 0.626 0.574 0.602 0.566 0.575 0.629 0.61 0.605 0.609 0.653
$F_{k}=0$ MASE 0.596 0.575 0.702 0.634 0.784 0.638 0.652 0.666 0.655 0.696 0.851 0.781
CRPS 0.533 0.513 0.653 0.572 0.707 0.568 0.58 0.623 0.601 0.628 0.756 0.735
@@
Dataset Metric TimesFM 2.5 Chronos-2 Kairos Moirai 2.0 VisionTS++ TiRex Toto Sundial TimesFM 2.0 Chronos-bolt Moirai TimesFM 1.0
Water Quality-Darwin MASE (raw) 0.993 0.938 1.220 1.092 1.270 1.024 1.007 1.292 1.182 1.036 1.428 1.098
@@
MASE (norm.) 0.739 0.706 0.928 0.816 0.989 0.766 0.754 0.953 0.888 0.779 1.072 0.827
CRPS (norm.) 0.504 0.466 0.694 0.584 0.795 0.553 0.500 0.745 0.645 0.563 0.675 0.630
MASE (rank) 2 1.333 10.667 6 10 4 3.333 11 9.333 4.667 12.333 6.667
@@
Current Velocity/5T MASE (raw) 0.729 0.710 0.956 0.929 1.183 0.789 0.833 0.907 0.804 1.008 1.281 0.897
@@
MASE (norm.) 0.601 0.588 0.787 0.752 1.005 0.647 0.678 0.738 0.662 0.819 1.046 0.734
CRPS (norm.) 0.649 0.632 0.786 0.757 0.862 0.666 0.691 0.749 0.712 0.780 0.903 0.766
MASE (rank) 2 1 9 8 11.333 3 4.667 7 4.333 10 12.333 6.667
@@
Current Velocity/10T MASE (raw) 0.695 0.686 0.815 0.734 0.884 0.703 0.738 0.768 0.738 0.746 0.904 0.814
@@
MASE (norm.) 0.648 0.640 0.754 0.682 0.840 0.655 0.685 0.714 0.685 0.693 0.836 0.752
CRPS (norm.) 0.659 0.655 0.697 0.675 0.708 0.658 0.677 0.710 0.689 0.677 0.716 0.717
MASE (rank) 2.333 1 10 5 10.667 2.667 5 8 5 7 11.667 9.667
CRPS (rank) 2.333 2 9 5.667 7.667 2 6.333 10 6.333 6 10.333 10.333
Current Velocity/15T MASE (raw) 0.762 0.746 0.866 0.860 1.017 0.779 0.814 0.897 0.796 1.017 1.094 0.864
CRPS (raw) 0.316 0.306 0.350 0.339 0.368 0.319 0.324 0.338 0.329 0.358 0.367 0.355
MASE (norm.) 0.644 0.631 0.736 0.723 0.898 0.656 0.687 0.758 0.673 0.852 0.927 0.726
CRPS (norm.) 0.648 0.627 0.716 0.693 0.760 0.655 0.665 0.695 0.673 0.734 0.754 0.722
MASE (rank) 2 1 7.333 6.667 9.667 3 5 9.333 4 11.333 12.333 7.667
CRPS (rank) 2 1 9 6.667 8.333 3.333 5 7.667 5 9.667 11 9.333
Current Velocity/20T MASE (raw) 0.567 0.603 0.790 0.787 0.876 0.639 0.665 0.621 0.608 0.743 0.848 0.719
CRPS (raw) 0.276 0.286 0.343 0.315 0.293 0.278 0.289 0.291 0.314 0.319 0.289 0.361
MASE (norm.) 0.545 0.573 0.733 0.715 0.875 0.608 0.624 0.596 0.583 0.699 0.811 0.675
CRPS (norm.) 0.641 0.661 0.769 0.720 0.683 0.645 0.670 0.678 0.720 0.727 0.672 0.815
MASE (rank) 1 3 10.333 9.667 10 4.667 4.667 4.333 3.333 9 11 7.667
CRPS (rank) 2 4 10 8 6.333 2 5.333 8 7 8.333 6.333 11.333
Current Velocity/H MASE (raw) 1.062 1.063 1.373 1.121 1.482 1.092 1.175 1.775 1.258 1.313 1.405 1.457
@@
MASE (norm.) 0.650 0.648 0.844 0.684 0.930 0.662 0.715 1.082 0.760 0.798 0.869 0.887
CRPS (norm.) 0.633 0.622 0.765 0.661 0.748 0.625 0.666 0.677 0.729 0.725 0.729 0.832
MASE (rank) 1.667 2 8.667 4 10.667 2.333 5 12.333 6.667 6.667 9 10
@@
CPHL/15T MASE (raw) 0.977 0.958 0.992 0.994 0.973 0.957 0.938 0.922 0.944 0.975 0.998 1.100
CRPS (raw) 0.213 0.212 0.225 0.216 0.213 0.213 0.208 0.216 0.212 0.218 0.222 0.252
MASE (norm.) 0.657 0.646 0.666 0.665 0.661 0.645 0.634 0.623 0.637 0.655 0.675 0.735
CRPS (norm.) 0.615 0.611 0.647 0.620 0.622 0.614 0.601 0.626 0.611 0.625 0.643 0.724
MASE (rank) 8 4.333 10 7.333 7.333 4.667 3.667 1.333 4 5.667 10 11.667
CRPS (rank) 4.667 4 10 5.333 5.333 5.667 3.333 6 4.667 7.667 9.333 12
CPHL/30T MASE (raw) 1.406 1.407 1.482 1.499 1.490 1.446 1.432 1.451 1.451 1.471 1.583 1.515
CRPS (raw) 0.248 0.250 0.270 0.267 0.266 0.262 0.255 0.287 0.262 0.268 0.285 0.286
MASE (norm.) 0.765 0.765 0.805 0.816 0.808 0.787 0.779 0.789 0.790 0.799 0.862 0.822
CRPS (norm.) 0.689 0.695 0.748 0.740 0.737 0.727 0.709 0.796 0.727 0.742 0.789 0.792
MASE (rank) 1.333 1.667 8.667 9.333 7.667 4.667 3 6.333 6.333 7 12 10
CRPS (rank) 1.333 1.667 8.333 7.333 6 4.667 3.333 11.333 5.667 6.667 10.667 11
CPHL/H MASE (raw) 1.645 1.679 1.699 1.699 1.828 1.714 1.748 1.841 1.705 1.744 1.801 1.869
CRPS (raw) 0.282 0.294 0.312 0.303 0.315 0.299 0.302 0.351 0.308 0.318 0.312 0.353
MASE (norm.) 0.800 0.812 0.825 0.824 0.890 0.829 0.841 0.888 0.827 0.847 0.875 0.900
CRPS (norm.) 0.743 0.770 0.815 0.795 0.829 0.783 0.791 0.916 0.805 0.833 0.820 0.918
MASE (rank) 1.667 2.667 4 4.667 10 5.333 6 10.333 5 7.333 9.667 11.333
CRPS (rank) 1 2 8 5.333 8.333 3.333 4.667 11.333 7 8.333 7.333 11.333
Coastal T-S/5T MASE (raw) 1.261 1.256 1.263 1.277 1.299 1.259 1.303 1.272 1.284 1.256 1.322 1.291
CRPS (raw) 0.072 0.068 0.081 0.070 0.076 0.068 0.069 0.093 0.080 0.075 0.069 0.082
MASE (norm.) 0.819 0.816 0.821 0.830 0.846 0.818 0.845 0.827 0.834 0.817 0.858 0.839
CRPS (norm.) 0.666 0.627 0.742 0.641 0.700 0.628 0.634 0.857 0.725 0.689 0.637 0.750
MASE (rank) 3 2.333 5 7 9 3 10.667 5.667 8.333 3.333 11.333 9.333
CRPS (rank) 6.667 1.667 10 4.667 9.333 2 2.667 11.333 7.667 8 4 10
Coastal T-S/15T MASE (raw) 0.881 0.958 1.271 1.131 1.276 1.013 1.102 1.488 1.268 1.579 2.107 1.074
CRPS (raw) 0.004 0.004 0.006 0.005 0.006 0.004 0.005 0.006 0.005 0.008 0.011 0.005
MASE (norm.) 0.547 0.608 0.809 0.715 0.878 0.621 0.673 0.873 0.762 0.948 1.258 0.674
CRPS (norm.) 0.512 0.571 0.877 0.708 0.995 0.593 0.710 0.860 0.717 1.082 1.456 0.697
MASE (rank) 1 2.667 9.333 6.667 7.333 3.333 4.333 10 8 10.667 12.333 5.333
CRPS (rank) 1 2.333 10 6.333 8.667 2.667 6 10 6.333 11 12.333 5.333
Coastal T-S/20T MASE (raw) 1.019 1.024 1.135 1.078 1.127 1.051 1.033 1.171 1.071 1.093 1.199 1.139
CRPS (raw) 0.043 0.042 0.055 0.043 0.046 0.042 0.044 0.057 0.054 0.048 0.045 0.056
MASE (norm.) 0.733 0.738 0.810 0.774 0.816 0.754 0.743 0.830 0.768 0.784 0.856 0.818
CRPS (norm.) 0.504 0.486 0.635 0.507 0.534 0.489 0.507 0.669 0.616 0.562 0.527 0.645
MASE (rank) 1.333 2 9.333 5.667 8.667 4 2.667 10.333 5.333 7.667 11.333 10